The expression x to the fourth -9 is equivalent to?
Answer by stanbon(75887)

The expression x to the fourth -9 is equivalent to?
----
x^4-9 = (x^2-3)(x^2+3) = (x-sqrt(3))(x+sqrt(3))(x^2+3)
----------------
Cheers,
Stan H.
--------------